### v1

@@ 3.4 Trading
A near universal difference between any CePM and DePM is that a DePM allows outcome shares to be withdrawn from the platform, typically in a form compliant with a token standard such as ERC-20 or the more efficient ERC-1155 [47]. Withdrawing outcome shares allows traders to exchange tokens outside of the platform and to compose with third party DeFi services, including on-chain trading, lending, and leverage [76]. Options for trading outcome shares can be broken into two steps: (i) how does the first outcome share come into existence and how does the first trader trade, and (ii) how do traders trade once a market has been established?
3.4.1 The first trade.
Probably the greatest evolution in DePMs, from Truthcoin to Polymarket, concerns how the first trade happens. There are three options: automated bookmaking, splitting, and matching.
Automated bookmaking was popularized through the academic work of Robin Hanson [33, 35], researched for Intrade [50], and first suggested for DePMs by Truthcoin [71], which heavily influenced Augur v1 [57]. In this model, the operator sets initial prices for each outcome share (equivalent to setting market odds) and collateralizes enough payout money to cover a worst-case loss in its treasury (preserving solvency in Definition 3). If Alice is the first trader, she can immediately trade with the operator. The operator is autonomous and sets buy/sell prices algorithmically, originally using Hanson’s logarithmic market scoring rule (LMSR) [35]. The key point is that the operator is Alice’s counterparty; if Alice wins, the operator loses, and vice-versa. The pros are instant liquidity for the first trader and the cons are the risk to the operator of losing money and the burden of needing to set initial odds (getting them wrong increases the chances it loses).
Acute readers might wonder the relationship between automated bookmaking and an automated market makers (e.g., Uniswap) discussed below. Roughly speaking, a WTA market run by automated bookmaking is termed a cost-function prediction market (CFPM) and a CFPM is equivalent in pricing (and trade costs) to an AMM (defined by a set of axioms) with the right invariant [27].
In YNB markets, each outcome’s YES/NO pair is its own two-outcome WTA market. Splitting is per outcome: converting \$1 of collateral mints one $j_{\mathsf{k_{Y}}}$ and one $j_{\mathsf{k_{N}}}$ for the chosen $\mathsf{k}$. Across outcomes, supplies are uncoupled—the total minted for the $\mathsf{A}$ bundle need not match that for $\mathsf{B}$ or $\mathsf{C}$.
The third approach, matching, was used by InTrade [37] and a variant by Fairlay [25]. Briefly, it mirrors a futures market, where Alice posts a desired short/long position at a chosen price on an orderbook with a margin account holding enough cash to cover her maximum loss if she obtains the position. If Bob is willing to take the other side, also with sufficient margin for his maximum loss, the operator matches them, creates two shares and gives them to Alice and Bob. Alice and Bob are not counterparties, both settle with the operator once the shares are created, however their coincidence of wants (COW) is necessary for the operator to create shares at no risk to itself. Matching is operationally more complex and effectively requires a central limit order book or batch auction, which excludes is too expensive to run on-chain [54].
3.4.2 Trading in established markets.
@@
Despite the direct lineage between prediction markets and AMMs, AMMs are problematic for prediction market trading. DePM outcome shares behave in specific ways that differ from typical crypto-assets and tokens. The price is strictly bounded between \$0 and \$1, the value of a share can jump to \$0 or \$1 near-instantly when an event outcome is finalized, and once finalized, the price is permanent. When real world events occur, AMMs can be drained faster than liquidity providers can withdraw liquidity. Adapting AMMs to these constraints is an interesting open problem. Paradigm’s pm-AMM tapers liquidity as a scheduled expiry approaches, which helps for events that crystallize over a fixed time horizon [53]; but many markets jump or resolve unexpectedly, falling outside of the model used for the results.
@@ 3.5 Market Resolution
It is possible that a market topic can be determined on-chain (e.g., total value locked in a DeFi service) in which case resolution is simple, however typically, prediction market outcomes concern facts that are off-chain. In these cases, resolution is the process of finalizing an off-chain outcome in an on-chain market. The predominate approaches are the following: self-settling markets, auto-resolve rules, using a designated arbiter, using a network of reporters, and crowdsourcing a vote. The term oracle is commonly used for any of the latter three approaches (while the first two could be considered oracle-less).
3.5.1 Oracle-less approaches.
A self-settling market is based splitting and merging (see §3.4) of shares and relies on participants with winning shares to purchase the losing shares (for close to \$0) and redeem \$1 by merging them [9]. If losing shares do not trade near \$0 or become illiquid, the broader market might accept winning shares as a substitute for dollars. If a market outcome is contentious with no recognized winner (e.g., a poorly defined market in §3.2), the market will not settle. An auto-resolve rule could also be used such that outcomes are finalized when, say, their token trades above \$0.99 for at least $t$ time. Auto-resolve rules are subject to market manipulation (losing shares are wash traded near \$1 long enough to finalize, in which case they are worth \$1) or griefing attacks (yes shares are traded below \$0.99 to prevent finalization). While an auto-resolve rule is too fragile to rely on solely, it could be used optimistically to provide an initial settlement; only disputed cases would be escalated to an oracle [3].
@@ 3.7 Archiving
4 Discussion and Research Agenda
Composability.
Table 3: Classification of selected (for significance or novelty) DePMs using our modular framework. Keywords used in the table are defined in referenced section, while ●, ◐, ○ respectively infer that the related property is fully, partially or not supported.
Regulation.
In jurisdictions like the United States, both CePMs and DePMs are tightly controlled and not always legal to operate [21, 49]. The paramount concerns of regulators are not specific to prediction markets or the difference in their microstructure from other forms of wagering (although market integrity is always a concern); rather, regulators control any kind of wagering on political events (a stance challenged by leading economists [2]), as well as any kind of financial service offered to unaccredited consumers. For advocates of liberal markets, DePMs are seen as a way to side-step regulation, particularly when the DePM is decentralized in most or all dimensions of our taxonomy.
Research Agenda.
The history of DePM shows a convergence toward trading outcome shares, creating by splitting, on CLOBs and AMMs. While many aspects of prediction markets have been studied formally, comparing different market microstructures has not received adequate research; and optimizing AMMs for DePMs is still at the beginning stages [42, 53]. Less convergence exists on how to decide market outcomes, a uniquely DePM problem (since a CePM can arbitrate its own markets). Another way to look for research problems is to pinpoint where faulty behaviour is still occurring: for Polymarket, the biggest sources are poorly defined market topics and the vulnerability of market resolution to manipulation. Analysis of new approaches to either problem (formal verification for topics? AI as oracles?) would be welcome research, as would bringing a theoretical lens to deployed oracle voting systems to understand if formal guarantees about incentives can be made. A lesser known issue is archiving DePM data and developing tools that could fully ‘replay’ a market at each timestep, which would be useful for understanding how markets incorporate news (real and fake) into prices. Finally, the degree to which DePMs can be pushed further toward permissionlessness in all aspects is still largely open, particularly around setting topics for markets.

### v2

@@ 3.5 Trading
Despite the direct lineage between prediction markets and AMMs, AMMs are problematic for prediction market trading. DePM outcome shares behave in specific ways that differ from typical crypto-assets and tokens. The price is strictly bounded between \$0 and \$1, the value of a share can jump to \$0 or \$1 near-instantly when an event outcome is finalized, and once finalized, the price is permanent. When real world events occur, AMMs can be drained faster than liquidity providers can withdraw liquidity.
Research Gap 7. Adapting AMMs to the unique constraints of predictions markets is at an early stage. Paradigm’s pm-AMM tapers liquidity as a scheduled expiry approaches, which helps for events that crystallize over a fixed time horizon [58]; but many markets jump or resolve unexpectedly, falling outside of the model used for the results.
@@
Research Gap 8. Can we measurement MEV extraction for on-chain DePMs? What mechanisms for order-flow privacy apply to on-chain DePMs?
@@ 3.6 Market Resolution
Market resolution approaches specify how a DePM determines and finalizes the winning outcome and triggers settlement. Designs vary in where the ground truth comes from (market prices, external sources, or participants), who can influence the decision, and what incentives/penalties secure it. In this section we review: oracle-less designs (on-chain truth, self-settling markets, auto-resolve rules), using a designated arbiter, using a network of reporters, and crowdsourcing a vote. The term oracle is commonly used for any of the latter three approaches.
3.6.1 Oracle-less.
A market topic might reference on-chain truth (e.g., total value locked in a DeFi service at the end of the year) in which case resolution can be made directly. A self-settling market [11] works with splitting and merging (see §3.5) of shares; it relies on participants with winning shares to purchase the losing shares (for close to \$0) and redeem \$1 by merging them. If losing shares do not trade near \$0 or become illiquid, the broader market might accept winning shares as a substitute for dollars. If a market outcome is contentious with no recognized winner (e.g., a poorly defined market in §3.2), the market will not settle. An auto-resolve rule could also be used such that outcomes are finalized when, say, their token trades above \$0.99 for at least $t$ time. Auto-resolve rules are subject to market manipulation (losing shares are wash traded near \$1 long enough to finalize, in which case they are worth \$1) or griefing attacks (yes shares are traded below \$0.99 to prevent finalization). While an auto-resolve rule is too fragile to rely on solely, it could be used optimistically to provide an initial settlement; only disputed cases would be escalated to an oracle [5].
@@
Research Gap 9. Deciding market outcomes is a uniquely DePM problem (since a CePM can arbitrate its own markets). Can we bring a theoretical lens to deployed oracle voting systems and understand what formal guarantees about incentives and safety can be made?
@@ 3.7 Settlement
Insight 5. A prediction market has no need to fight against paying out a bet the way bookmakers have in practice [31]. When operations are on-chain, solvency can be demonstrated directly from blockchain data.
@@ 3.8 Archiving
Research Gap 10. A useful tool that does not exist yet could use on-chain data, API data, and external data (news, social media, etc.) to replay a past market, allowing researchers to step forward and backward at each timestamp. It was onerous for us to manually study how how markets incorporate news (real and fake) into prices in §0.B without such a tool.